OpenStudy (anonymous):

Find the equation of the line. the line passthrough (8, 4) and has the same x intercept as the line -4x+y=1 write in form y=mx+b

Directrix (directrix):

Find the x-intercept of -4x+y=1. Let y = 0. -4x + 0 = 1 -4x = 1 x = - 1/4 Now to find the equation of the line joining (-1/4,0) and (8,4) ....

Directrix (directrix):

The slope of the line is [ 0 - 4] / (-1/4) - 8] = (-4) / (-33/4) = 16/33. Using the point (8,4), y - 4 = (16/33) (x - 8)

Directrix (directrix):

Multiply through by 33. 33y - 132 = 16(x - 8)

OpenStudy (anonymous):

is it -4

Directrix (directrix):

33y - 132 = 16x - 128

Directrix (directrix):

33y = 16x + 4

Directrix (directrix):

y = (16/33) x + 4/33

Directrix (directrix):

That is the equation -- the answer.

OpenStudy (anonymous):

THANK YOU SO MUCH YOUR AWSOME

Directrix (directrix):

Glad to help.